# PRISM: Prompt-Refined In-Context System Modelling for Financial Retrieval

### v1

@@ 5 Experiments
We evaluated four GPT models (GPT-4o-mini, GPT-4.1, GPT-5-mini, and GPT-5) and employed OpenAI’s text-embedding-3-small (TE3-S) and text-embedding-3-large (TE3-L) for ICL retrieval. Performance is measured using NDCG@5, with results reported on a validation split of 30% public and 70% private subsets. On the private subset, PRISM achieved performance close to the top-performing models, with an average gap of 0.006 and a margin of over 0.02 from the remaining ones. Exact methods and scores are omitted to preserve anonymity under leaderboard-based evaluation.
5.1 Ablation Studies
@@
System Prompt Engineering. The complete ablation results are shown in Table 3. We first analyze the effects of prompt design and model capacity in the non-agentic workflow. Early prompt variants ($P_{1}$$P_{2}$) revealed that concise, direct instructions improved retrieval relevance over overly structured formats. Later versions ($P_{3}$$P_{4}$) emphasized clarity, reasoning focus, and contextual grounding, further stabilizing performance. The $P_{4}$ prompt, which encouraged explicit reasoning about document relevance while reducing verbosity, produced steady gains and improved the NDCG@5 from 0.64297 (Run 12) to 0.65559 (Run 13) with GPT-5-mini. Increasing model capacity from GPT-4.1 to GPT-5-mini and GPT-5 consistently enhanced performance, showing that larger models follow structured, reasoning-oriented prompts more effectively. Run 15 has the best results among all variants, with systematic reasoning guided by our prompt design to rank relevant documents and chunks accurately.
@@
In-Context Learning. The second study examines the impact of ICL on both non-agentic and agentic workflows. In the non-agentic workflow, applying ICL at the document level consistently improved performance, while extending it to both document and chunk levels degraded results due to context overload and fragmented attention. The best configuration applied ICL only at the document stage, achieving the highest score (Run 19, 0.71818), demonstrating that strategically scoped ICL enhances reasoning consistency, whereas excessive context dilutes retrieval focus. In contrast, the agentic workflow showed mixed results. Increasing the number of shots often reduced performance (Run 28 vs. Runs 29–32), indicating that excessive or poorly selected examples can bias the model’s internal policy and impair generalization. A modest setup, such as 5-shot ICL with GPT-5-mini (Run 28), provided balanced gains and show that limited examples can guide reasoning without over constraining it. Overall, ICL is most effective when applied selectively as structured guidance to improve alignment in financial IR tasks.
@@
Multi-Agent Workflows. The third study examines the performance dynamics of MAS under different model scales and graph configurations ($A_{1}$$A_{4}$). Each design varies in the number and connectivity of agents handling filtering, ranking, and scoring. Results show that MAS performance is highly sensitive to model size and architectural complexity. Smaller models, such as GPT-4o-mini, struggled in deeper graphs like $A_{2}$ (Run 22, 0.51518) due to error propagation across agents. In contrast, scaling to GPT-5-mini significantly improved stability and accuracy, with the $A_{3}$ workflow rising from 0.53400 (Run 24) to 0.66291 (Run 25). This highlights that larger models better sustain coherent multi-agent coordination across complex reasoning chains. Additional ablations further reveal that hybrid configurations with agentic at the document level but non-agentic at the chunk level achieved strong results (0.70685, Run 33; 0.69439, Run 34), comparable to the best overall performance. This indicates that document level agentic reasoning is effective in isolation, whereas extending agentic control to chunk ranking introduces excessive coordination overhead. More findings are reported in Limitations and Appendix.
@@ 6 Feasibility and Reproducibility Analysis
We report the statistical evaluation of model performance across repeated runs in this section.
Feasibility Analysis. Runs of each prompt variant and ICL are selected to evaluate the practical feasibility of PRISM by analysing latency and token efficiency. Figure 3 shows that document ranking tasks have low median latency (8–10s) with minimal variance, while chunk ranking tasks are slower (130–160s) due to higher reasoning complexity. Runs 15 and 19 strike the best balance between contextual depth and stability. As shown in Figure 4, document-ranking tasks average 2–3K tokens per sample (70% prompt, 30% completion), whereas chunk-ranking tasks use around 100K tokens but remain prompt-dominant (8̃5%). In short, Runs 15 and 19 deliver consistent performance with predictable latency and efficient token usage. Detailed cost analyses are also provided in the Appendix.
Figure 3: Latency distribution box plots.
Figure 4: Token distribution bar charts.
Reproducibility Analysis. Run 12 serves as the baseline, with Runs 15, 18, and 19 using improved prompts and larger models. All runs exhibit low variability, with standard deviation ($s<0.011$) and coefficient of variation (CV $1.6\%$), indicating stable and reproducible outcomes. The narrow 95% confidence intervals (CI) confirm the statistical reliability of the mean performance estimates. Run 12 recorded the lowest mean score, while Runs 15–19 achieved consistently higher averages, with non-overlapping CIs showing statistically significant gains following configuration updates. Welch’s two-sample $t$-tests welch1938significance further support these improvements ($p<0.05$ for all Run 12 comparisons). In general, the low variance and narrow confidence bounds demonstrate that the updated configurations produce statistically significant and reproducible results.
@@
Table 5: Welch’s $t$-tests of Run 12 vs. subsequent runs.
Comparison $t$-statistic $p$-value Significance ($\alpha=0.05$)
Run 12 vs Run 15 -3.96943 0.00573 Significant
Run 12 vs Run 18 -4.98117 0.00218 Significant
Run 12 vs Run 19 -6.58157 0.00141 Significant
@@ 7 Conclusion
This work presents PRISM, a training-free framework that integrates system prompting, in-context learning, and multi-agent system modelling. Extensive ablation studies show that non-agentic workflows using the $P_{4}$ prompt and document-level ICL achieved outstanding performance on FinAgentBench. In contrast, multi-agent configurations faced overfitting and coordination overhead. These results demonstrate that PRISM is both effective and reproducible, making it a practical solution for production-scale financial information retrieval.
Limitations
Data Preprocessing
Non-Agentic Workflow
Although the non-agentic configuration achieved the strongest overall performance, retrieval fidelity remains a notable limitation. Our best ICL setup relies heavily on the quality of the TE3-S model, and currently uses a simple L2 distance–based embedding lookup to retrieve relevant chunks. More advanced retrieval strategies could yield higher-quality candidates. A future improvement could involve hybrid retrieval that integrates dense semantic embeddings with lexical methods such as BM25 bm25, combined through Reciprocal Rank Fusion reciprocalrankfusion. This would provide a more balanced retrieval signal, capturing both semantic context and domain-specific financial terminology that purely dense methods may overlook.
Agentic Workflow
Our agentic experiments revealed two main challenges. First, multi-agent systems are highly sensitive to prompt–architecture compatibility: the $P_{4}$ prompt, which consistently improved non-agentic performance, led to a substantial drop (0.55757) in MAS. This suggests that constraint-heavy prompts can disrupt inter-agent communication. Second, tuning graph depth and coordination remains difficult as deeper or more interconnected topologies often introduce instability rather than improving reasoning. These findings highlight the need for lighter prompts and adaptive coordination mechanisms tailored for distributed reasoning.
Architectural Design
Architecturally, our results indicate a saturation bottleneck in the current MAS design. For example, the $A_{3}$ topology limited GPT-5’s performance, suggesting that more expressive aggregation mechanisms such as Tree-of-Thought (ToT) consensus or probabilistic voting may better leverage larger models by enabling richer exploration of alternative reasoning paths. This can potentially break through the saturation point we observed and allowing superior models to fully leverage MAS.
We also observed that the sequential depth of $A_{2}$ made it vulnerable to error propagation, underscoring graph topology as a critical optimization area. Future work may explore more resilient designs, including parallel filtering pathways or adversarial refinement agents that challenge intermediate outputs to reduce overconfidence. Such mechanisms, combined with more parallel rather than sequential structures, could reduce compounding errors and improve the scalability of agentic workflows.
@@ Appendix A Appendix
In this Appendix we provide additional information and insights that cannot be fit into main paper due to the page limit. We follow the sequence of sections in the main paper for easier understanding.
A.1 Chunk Ranking Data Analysis
@@
We conducted a frequency analysis on chunk word counts to complement the token count analysis and validate our observations. As shown in Figure 5, the trends closely mirror the token count results in Figure 2 where the relevant chunks consistently contain more words, reflecting their higher informational density and richer semantic content. The similar distribution patterns, including wider interquartile ranges and higher upper extremes, further corroborate the variability observed in chunk lengths. This strong alignment between word and token count analyses reinforces the reliability of our insights. Adaptive retrieval and ranking strategy is needed to identify detailed, content-heavy chunks with mechanisms to handle large fluctuations in chunk length efficiently.
A.2 Dynamics of In-Context Learning and Prompt Versioning

### v2

@@ 5 Experiments
We evaluated four GPT models (GPT-4o-mini, GPT-4.1, GPT-5-mini, and GPT-5) using OpenAI’s text-embedding-3-small (TE3-S) and text-embedding-3-large (TE3-L) for ICL retrieval. Our evaluation spans three datasets selected to cover complementary task characteristics: FinAgentBench (document and chunk ranking with SEC filings), FiQA-2018 (passage reranking for financial QA), and FinanceBench (factual QA requiring numerical reasoning). While our experiments use OpenAI models exclusively due to infrastructure constraints (see Appendix A.8), the training-free nature of PRISM allows straightforward adaptation to other LLM providers.
5.1 Quantitative Results
FinAgentBench. Performance is measured using NDCG@5, with results reported on a validation split of 30% public and 70% private subsets. As ground truth labels for the public and private test sets are not provided, we report only the combined document and chunk ranking scores computed by the Kaggle evaluation system. Following the competition’s conclusion, this evaluation system became defunct, preventing further experiments. On the private subset, PRISM ranked 3rd overall while being the only training-free method in the top three. The top two methods employed fine-tuned embedding models (30 epochs and three separate MiniLM models respectively), whereas PRISM achieves competitive performance using only off-the-shelf LLMs with prompt engineering and ICL. Our best observed score across repeated runs of the same configuration (Run 19) is 0.71818, trailing the top method by only 0.007 (0.72497 vs. 0.71818) while leading the remaining methods by approximately 0.02, indicating that training-free approaches can be competitive with fine-tuned systems when properly configured.
Table 4: FinanceBench oracle evaluation results.
Prompt Model ICL Correct Incorrect Unable to Answer
Baseline GPT-5 N/A 141 9 0
$P_{1}$ GPT-5 ICL-9 147 3 0
$P_{2}$ GPT-5 ICL-9 142 8 0
$P_{3}$ GPT-5 ICL-9 140 10 0
$P_{4}$ GPT-5 ICL-9 146 4 0
Table 5: FiQA evaluation results.
Prompt Model ICL NDCG@10 Recall@100
BM25+CE N/A N/A 0.3470 0.5390
$P_{1}$ GPT-5 ICL-5 0.6104 0.8083
$P_{2}$ GPT-5 ICL-5 0.6027 0.7941
$P_{3}$ GPT-5 ICL-5 0.6197 0.8206
$P_{4}$ GPT-5 ICL-5 0.6097 0.8021
FiQA-2018. As shown in Table 5, PRISM achieves substantial improvements over existing methods on the test set, with gains of 78.59% in NDCG@10 and 52.24% in Recall@100 compared to the best-performing reranking baseline, BM25+Cross-Encoder (CE). Baselines from the BEIR benchmark (Thakur et al., 2021) are reported directly from published results under the standard BEIR evaluation protocol. With GPT-5, PRISM consistently achieves NDCG@10 exceeding 0.60 and Recall@100 reaching close to 0.80 across all prompt variations, indicating robust performance across different configurations. All PRISM configurations outperform the strongest traditional baseline (BM25+CE). Comprehensive model comparisons against all BEIR baselines and feasibility analysis are provided in Appendix Table 10 and Table LABEL:tab:full_fiqa_results.
FinanceBench. PRISM is also evaluated on FinanceBench under oracle context conditions (Table 5), replicating the best-performing baseline from the original paper with a similar model. PRISM achieved 98% accuracy with the $P_{1}$ prompt configuration, surpassing the baseline. Notably, while prompt variations performed comparably on reranking datasets, the task-specific nature of QA led to $P_{1}$ slightly outperforming other variants, indicating that optimal prompt formulations are task-dependent. All PRISM configurations exceeded 90% accuracy, demonstrating adaptability across retrieval paradigms. Complete results are provided in Appendix A.9.3 and Table LABEL:tab:full_financebench_results.
5.2 Ablation Studies on FinAgentBench
@@
System Prompt Engineering. Based on the results in Table 3, we first analyze the effects of prompt design and model capacity in the non-agentic workflow. Early prompt variants ($P_{1}$$P_{2}$) revealed that concise, direct instructions improved retrieval relevance over overly structured formats. Later versions ($P_{3}$$P_{4}$) emphasized clarity, reasoning focus, and contextual grounding, further stabilizing performance. The $P_{4}$ prompt, which encouraged explicit reasoning about document relevance while reducing verbosity, produced steady gains and improved the public NDCG@5 from 0.64297 (Run 12) to 0.65559 (Run 13) with GPT-5-mini. Increasing model capacity from GPT-4.1 to GPT-5-mini and GPT-5 enhanced performance, showing that larger models follow structured, reasoning-oriented prompts more effectively. Run 15 achieves the best results among these configurations, with systematic reasoning guided by our prompt design to rank documents and chunks accurately.
@@
In-Context Learning. The second study examines the impact of ICL on both non-agentic and agentic workflows. In the non-agentic workflow, applying ICL at the document level consistently improved performance, while extending it to both document and chunk levels degraded results due to context overload and fragmented attention. The best configuration applied ICL only at the document stage, achieving the highest score (Run 19, 0.71818), demonstrating that strategically scoped ICL enhances reasoning consistency, whereas excessive context dilutes retrieval focus. In contrast, the agentic workflow showed mixed results. Increasing the number of shots often reduced performance (Run 28 vs. Runs 29–32), indicating that excessive or poorly selected examples can bias the model’s internal policy and impair generalization. A modest setup, such as ICL-5 with GPT-5-mini (Run 28), provided balanced gains and shows that limited examples can guide reasoning without over-constraining it. Overall, ICL is most effective when applied selectively as structured guidance to improve alignment in financial IR tasks.
@@
Multi-Agent Workflows. The third study examines the performance dynamics of MAS under different model scales and graph configurations ($A_{1}$$A_{4}$). Each design varies in the number and connectivity of agents handling filtering, ranking, and scoring. Results show that MAS performance is highly sensitive to model size and architectural complexity. Smaller models such as GPT-4o-mini struggled in deeper graphs like $A_{2}$ (Run 22, 0.51518) due to error propagation across agents, whereas scaling to GPT-5-mini significantly improved stability and accuracy, with the $A_{3}$ workflow rising from 0.53400 (Run 24) to 0.66291 (Run 25). This model-size dependency is a key empirical finding that MAS benefits emerge only when individual agents possess sufficient reasoning capacity to maintain coherent intermediate outputs while coordination overhead outweighs the benefits of task decomposition for smaller models. Additional ablations reveal that hybrid configurations of agentic at the document level but non-agentic at the chunk level achieved strong results (0.70685, Run 33; 0.69439, Run 34), comparable to the best overall performance. This indicates that document-level agentic reasoning is effective in isolation, whereas extending agentic control to chunk ranking introduces excessive coordination overhead. Findings from these failure modes are discussed further in Appendix A.13 and Appendix A.7.
@@ 6 Feasibility and Reproducibility Analysis
While prior work on FinAgentBench focuses exclusively on ranking accuracy, practical deployment requires understanding computational costs. To our knowledge, we provide the first latency and cost analysis for this benchmark, which also establishes a baseline for evaluating accuracy-latency-cost trade-offs in future systems. Run 19 achieves average latencies of 10.43s for document ranking and 131.13s for chunk ranking, with per-query token costs of \$0.0155 and \$0.2768 respectively. These metrics meet typical enterprise search latency requirements (sub-minute for document retrieval) while the cost structure ($\sim$\$0.30 per complex query) remains viable for high-value financial analysis where accuracy outweighs inference costs. Limitations of PRISM are discussed in Appendix A.13.
Figure 3: Latency distribution box plots, where each box shows how the latency varies across different runs that are affected by different LLMs and configurations. For chunk ranking tasks, we observe that $P_{3}$ (Run 12) has the highest median latency for both tasks due to its heavy reasoning prompt design, while $P_{4}$ (Runs 13–19) achieves a good balance between latency and ranking performance.
Feasibility Analysis. Runs of each prompt variant and ICL are selected to evaluate the practical feasibility of PRISM by analyzing latency and token efficiency. Figure 3 shows that document ranking tasks have low median latency (3–17s) with minimal variance, while chunk ranking tasks are slower (32–156s) due to higher reasoning complexity. Runs 15 and 19 strike the best balance between contextual depth and stability. As shown in Figure 4, document-ranking tasks average 2–3K tokens per sample (70% prompt, 30% completion), whereas chunk-ranking tasks use around 100K tokens but remain prompt-dominant (85%). In short, Runs 15 and 19 deliver consistent performance with predictable latency and efficient token usage. Detailed cost analyses and a production readiness evaluation are provided in Appendix A.10.1 and Appendix A.12.
Figure 4: Token distribution bar charts, where each bar details the proportion of prompt and completion token usage across different runs and tasks, highlighting that chunk ranking tasks use more prompt tokens due to longer context and complexity.
Reproducibility Analysis. Run 12 serves as the baseline, with Runs 15, 18, and 19 using improved prompts and larger models. As detailed in Tables 7 and 7, all runs exhibit low variability (CV $1.6\%$), and Welch’s two-sample $t$-tests (Welch, 1938) confirm statistically significant improvements for all comparisons against Run 12 ($p<0.05$). Further discussion is provided in Appendix A.11.
@@
Table 7: Welch’s $t$-tests of Run 12.
Comparison $t$-stat $p$-val Signif.
12 vs 15 -3.969 0.0057 Yes
12 vs 18 -4.981 0.0022 Yes
12 vs 19 -6.582 0.0014 Yes
@@ 7 Conclusion
This work presents PRISM, a training-free framework that integrates system prompting, in-context learning, and multi-agent coordination. Rather than claiming methodological novelty, our primary contribution is a systematic empirical study of when each component provides value in financial retrieval. Our extensive ablation studies reveal that simpler non-agentic workflows using the $P_{4}$ prompt with document-level ICL achieved the strongest overall performance (NDCG@5 of 0.71818), while multi-agent configurations faced coordination overhead and model-size dependencies that limited their effectiveness, challenging common assumptions about the benefits of agentic architectures. Importantly, our study provides practical guidance: prompt engineering offers reliable performance with minimal overhead, ICL enhances reasoning when applied selectively at the document level, and MAS shows potential only with larger models and careful scope limitation. Evaluation across three datasets (FinAgentBench, FiQA-2018, FinanceBench) demonstrates PRISM’s cross-dataset adaptability, while our feasibility analysis provides the first latency-cost baselines for training-free financial IR. Extending this work to open-source model families remains an important direction for future work.
8 Ethical Use of Data
All datasets used in this work are publicly available and were obtained through their official distribution channels. FinAgentBench (Choi et al., 2025b) was released as part of an open Kaggle competition and contains anonymized SEC filing excerpts intended for research use. FiQA-2018 is distributed through the BEIR benchmark (Thakur et al., 2021) as a standard open-access evaluation resource. FinanceBench (Islam et al., 2023) is publicly released for non-commercial research purposes and consists of questions grounded in publicly filed financial documents. No proprietary data, personally identifiable information, or sensitive records were collected, stored, or used at any stage of this work. The datasets involve financial text drawn from public regulatory filings, which carry no privacy implications. Our use of these resources is fully consistent with their intended research purposes, and this work raises no ethical concerns.
9 Reproducibility
PRISM is training-free and introduces no learned parameters, meaning reproduction requires no model training or fine-tuning. All experiments were conducted using fixed, versioned OpenAI models (GPT-4o-mini, GPT-4.1, GPT-5-mini, GPT-5) with exact model identifiers documented in Appendix A.8. Prompt templates, ICL retrieval configurations, MAS architectural variants, and all hyperparameters are fully specified therein. To assess run-to-run consistency, we repeated key configurations across multiple independent trials; the resulting descriptive statistics, confidence intervals, and Welch’s $t$-test results are reported in Appendix A.11, confirming low variance (CV $1.6\%$) across runs. The full source code, including prompt files and pipeline configurations, is publicly available at https://bit.ly/prism-ailens.
10 Acknowledgement
We thank the Cradle Fund and the Microsoft for Startups program for their funding support, which made this research possible. We also thank the DocLab, DocSuite, and MLSuite teams at AI Lens for their contributions which collectively shaped the PRISM pipeline presented in this work.
@@ Appendix A Appendix
In this appendix, we provide additional information and insights that cannot fit into the main paper due to the page limit. We follow the sequence of sections in the main paper for easier understanding.
A.1 Exploratory data analysis (EDA) on FinAgentBench
A.1.1 Document Ranking Keyword Analysis
Table 2 reveals distinct keyword concentration patterns across SEC filing types. DEF 14A filings are dominated by high-ratio keywords such as dependency (88.40%) and concentration (88.24%), while 8-K filings focus more on compensation-related terms (95.45%). 10-Q filings contain notably lower keyword ratios, reflecting greater variation in language use. In contrast, Earnings Transcripts and DEF 14A documents show stronger keyword concentration, suggesting that certain document types have more distinctive language patterns that can help guide retrieval. These patterns informed the design of our domain-specific prompt priors and agent routing logic.
A.1.2 Chunk Ranking Data Analysis with Word Count
@@
We conducted a frequency analysis on chunk word counts to complement the token count analysis and validate our observations. As shown in Figure 5, the trends closely mirror the token count results in Figure 2 where the relevant chunks consistently contain more words, reflecting their higher informational density and richer semantic content. The similar distribution patterns, including wider interquartile ranges and higher upper extremes, further corroborate the variability observed in chunk lengths. This strong alignment between word and token count analyses reinforces the reliability of our insights. An adaptive retrieval and ranking strategy is needed to identify detailed, content-heavy chunks, with mechanisms to handle large fluctuations in chunk length efficiently.
A.2 EDA on FiQA-2018
Figure 6: Distribution of relevant documents per query.
FiQA-2018 contains 57,638 documents and 648 queries with 1,706 relevance judgments. With only 2.96% of the corpus marked as relevant and a dataset sparsity of 99.9954%, FiQA-2018 exhibits extreme sparsity typical of real-world retrieval scenarios where relevant information constitutes a small fraction of the available corpus. Figure 6 shows the distribution of relevant documents per query. The majority of queries have fewer than 3 relevant documents, with a median of 2 and a mean of 2.63 relevant documents per query. Despite this concentration, the distribution has a long tail extending up to 15 relevant documents for certain queries, indicating significant variability in query complexity. The IQR between Q1 and Q3 is narrow, confirming that most queries target specific, focused information. This fine-grained relevance pattern, combined with the extremely sparse judgment set, requires retrieval models to discriminate precisely between relevant and irrelevant content within a vast corpus of financial documents. The relatively short average query length of 11 words and document length of 133 words suggest that both queries and documents are concise, placing additional emphasis on semantic understanding rather than keyword matching.
A.3 EDA on FinanceBench
Table 8: Top 6 keywords in each document type; ratios show corpus-relative frequency.
10-K 10-Q 8-K Earnings
Keyword Ratio Keyword Ratio Keyword Ratio Keyword Ratio
securely 99.00% mun 98.61% accession 96.55% twentieth 58.90%
restatement 99.00% min 85.71% chargeable 80.00% fragrance 43.33%
athletic 99.00% noncompensation 83.13% survivorship 77.27% recalculate 42.31%
george 98.99% aum 80.95% abstain 77.27% translational 40.24%
budgetary 98.99% logged 80.77% unbudgeted 74.36% eighteenth 40.00%
sport 98.99% nettable 80.00% bailiwick 73.85% shasta 36.21%
FinanceBench contains 150 documents across four types: 10-K reports (112), 10-Q reports (15), 8-K reports (9), and Earnings Transcripts (14), spanning 32 unique companies. Table 8 presents the top 6 keywords for each document type. The analysis reveals distinct lexical patterns across document types. 10-K reports exhibit highly concentrated company and industry-specific vocabulary (98-99%) with terms like ’athletic’ and ’sport’ reflecting sector specialization. 10-Q and 8-K filings show strong concentration (74-98%) with specialized financial and procedural terminology. Earnings transcripts demonstrate the most diverse vocabulary (36-58%) with comparative terms like ’twentieth’ and ’recalculate’, suggesting broader conversational patterns. These distinct vocabularies suggest that developing retrieval strategies to specific document types could improve performance.
A.4 Different Prompt Versions
The non-agentic prompt templates are designed for single LLM workflows, where the model independently performs the complete ranking process using structured system instructions. These prompts emphasize logical consistency, domain alignment, and interpretable reasoning. For document ranking, the progression moves from domain-guided heuristics to combined reasoning frameworks. For chunk ranking, prompts evolve from a pure ReAct framework to integrated multi-strategy approaches. Specifically, the prompt iterations are as follows:
  • $P_{1}$: For document ranking, introduces domain-specific keyword hints and statistical cues to guide relevance assessment. For chunk ranking, employs a ReAct-style framework that interleaves internal reasoning with action steps, requiring the model to think step-by-step before outputting ranked indices.
  • $P_{2}$: For document ranking, enhances $P_{1}$ with detailed statistical keyword associations (e.g., word frequency ratios per document type) to improve discriminative power. For chunk ranking, combines Chain-of-Thought (CoT), ReAct, and Tree-of-Thought (ToT) strategies, introducing multi-expert simulation where three analysts reason collaboratively and prune incorrect paths.
  • $P_{3}$: Adopts a unified multi-strategy framework for both document and chunk ranking, integrating CoT decomposition, ReAct alternation, and ToT multi-expert simulation. Emphasizes explicit step-by-step reasoning with domain-specific hints for financial concepts.
  • $P_{4}$: Refines the $P_{3}$ framework with streamlined instructions and consistent terminology across document and chunk ranking tasks, maintaining the combined CoT, ReAct, and ToT reasoning strategies.
The complete prompt templates for document ranking (Table LABEL:tab:doc-prompts) and chunk ranking (Table LABEL:tab:chunk-prompts) are provided in Appendix A.14.
A.5 Dynamics of In-Context Learning and Prompt Versioning